Find the general solution

1) x3y'''-4x2y''+8xy'-8y=4lnx


1
Expert's answer
2021-08-05T17:19:41-0400

Let us solve the differential equation "x^3y'''-4x^2y''+8xy'-8y=4\ln x." Let us use the transformation "x=e^t." Then "y'_x=y'_t\cdot t'_x=y'_t\cdot \frac{1}{x'_t}=y'_t\cdot e^{-t}," "y''_{x^2}=(y''_{t^2}\cdot e^{-t}-y'_t\cdot e^{-t})e^{-t}=(y''_{t^2}-y'_t)e^{-2t}," "y'''_{x^3}=(y'''_{t^3}-3y''_{t^2}+2y'_t)e^{-3t}."

It follows that we have the following equation "e^{3t}(y'''_{t^3}-3y''_{t^2}+2y'_t)e^{-3t}-4e^{2t}(y''_{t^2}-y'_t)e^{-2t}+8e^ty'_te^{-t}-8y=4t," which is equivalent to the equation "y'''_{t^3}-7y''_{t^2}+14y'_t-8y=4t." The characteristic equation "k^3-7k^2+14k-8=0" is equivalent to "(k-1)(k^2-6k+8)=0," and to "(k-1)(k-2)(k-4)=0," and hence has the roots "k_1=1, k_2=2" and "k_3=4." The general solution of the differential equation "y'''_{t^3}-7y''_{t^2}+14y'_t-8y=4t" is "y(t)=C_1e^t+C_2e^{2t}+C_3e^{4t}+y_p(t)," where "y_p(t)=at+b.\\" It follows that "y_p'(t)=a,y_p''(t)=0,y_p'''(t)=0." We have that "14a-8(at+b)=4t." Then "-8a=4, 14a-8b=0," and thus "a=-\frac{1}{2}, b=\frac{7}{4}a=-\frac{7}{8}." Therefore, "y(t)=C_1e^t+C_2e^{2t}+C_3e^{4t}-\frac{1}{2}t-\frac{7}{8}.\\" We conclude that the general solutions of the differential equation "x^3y'''-4x^2y''+8xy'-8y=4\ln x" is "y(x)=C_1x+C_2x^2+C_3x^4-\frac{1}{2}\ln x-\frac{7}{8}."